# Speculative Safety-Aware Decoding

### v1

@@ Abstract
Xuekang Wang1  Shengyu Zhu2thanks: Corresponding author: zhushengyu@ict.ac.cn  Xueqi Cheng2 1 School of Computer Science & Technology, Beijing Institute of Technology 2 State Key Laboratory of AI Safety, Institute of Computing Technology, CAS wangxk@bit.edu.cn  {zhushengyu,cxq}@ict.ac.cn
@@ 5 Experiment
This section evaluates the effectiveness, helpfulness and efficiency of the proposed method SSD.
@@ 5.1 Experiment Setup
We evaluate SSD on three open-source large models, namely Vicuna-7b Chiang et al. (2023), Llama2-7b-chat and Llama2-13b-chat Touvron et al. (2023), as target models, and a small model TinyLlama-1.1B-Chat (Zhang et al., 2024a) as the expert model. Notice that only TinyLlama is fine-tuned to possess the deep safety alignment property following Qi et al. (2025).
@@
For utility, we consider two datasets: GSM8K (Cobbe et al., 2021) and Just-Eval (Lin et al., 2023). GSM8K is designed to test the model’s ability to solve complex mathematical problems. Notice that TinyLLama demonstrates very low accuracy on GSM8K and this task would challenging for the composite model. We use standard ROUGE-1 score for the accuracy on GSM8K, in line with existing evaluation practices. Similar to Xu et al. (2024), we use 1000 diverse instructions from Just-Eval to evaluate the outputs in terms of helpfulness, clarity, factuality, depth, and engagement.
@@ 5.2 Experimental Results
Table˜3 reports the Average Token Generation Time Ratio (ATGR), comparing the inference speed of our method, SSD, against other defense strategies like Deep-Align and SafeDecoding. The results validate that SSD can successfully accelerate decoding ($\text{ATGR}<1$), particularly for larger models. In contrast, SafeDecoding’s ATGR is greater than $1$ because its design of safety checks adds computational overhead and cannot utilize the auxiliary model’s outputs to accelerate inference. It is also worth noting that the ATGRs does not account for the cost of fine-tuning, which demands far greater computational resources for large models.
@@
SSD Avoids Over-refusal Responses on Borderline Queries.
As shown in Table˜4, SSD demonstrates a strong ability to handle borderline cases. Compared to the No Defense baseline, SSD introduces only a minimal increase in false refusals, indicating that its safety enhancements are well-targeted. In contrast, other defense methods exhibit a significantly higher FRR, suggesting they are prone to over-refusal behavior. This highlights that SSDeffectively defends against genuine threats while maintaining helpfulness on nuanced, sensitive topics.
Method False Refusal Rate $\downarrow$
LLM as judge String matching
No Defense 20.00% 26.80%
Deep-Align 29.60% 49.20%
SafeDecoding 28.40% 65.60%
SSD (Ours) 20.80% 30.40%
@@ 7 Limitations
In this paper, the largest model on which we evaluate the proposed method is Llama2-13b. Due to limited GPU resources, we cannot conduct experiments to validate how our method applies to larger models like Llama2-70B. Another limitation of the current work is on the difference of conversational style between the original and the expert models, e.g., Vicuna and TinyLlama, which we do not further investigate. Future research can take into account this difference of conversational style to improve performance w.r.t. both model utility and safety for more practical scenarios.
@@ 8 Ethics Impact
This paper develops a lightweight and efficient method to strengthen existing LLMs with additional safety properties to defend against new jailbreak attacks. We empirically show that the developed method can effectively and efficiently equip the LLM with the deep safety alignment property and further improve the original model’s capacity against many types of jailbreaks. This research aims to enhance the safety of large models and to contribute positively to the broader field of AI research. We remark that the development of SSD only uses publicly available jailbreak prompts and do not create new ones. In the paper, only one jailbreak input query is exhibited in an abstracted way, for illustration purpose. We will also release the codes to facilitate red-teaming efforts on LLMs.
@@ B.4 Setup of Deep-Align
We train the Llama2-7b and Vicuna models with deep safety alignment properties using the same hyperparameters and datasets as the default settings in Qi et al. (2025). For the TinyLlama used as the small expert model $m$, we adjust the learning rate of the hyperparameters to $2\times 10^{-4}$.

### v2

@@ Abstract
Xuekang Wang2   Shengyu Zhu1thanks: Corresponding author: zhushengyu@ict.ac.cn   Xueqi Cheng1 1 State Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2 School of Computer Science & Technology, Beijing Institute of Technology wangxk@bit.edu.cn  {zhushengyu,cxq}@ict.ac.cn
@@ 5 Experiment
This section evaluates the effectiveness, helpfulness and efficiency of the proposed method SSD. An implementation of SSD has been made available at https://github.com/k-k1w-w1x-x/Speculative-Safety-Aware-Decoding.
@@ 5.1 Experiment Setup
We evaluate SSD using open-source large models, namely, Vicuna-7b Chiang et al. (2023), Llama2-7b-chat and Llama2-13b-chat Touvron et al. (2023), as target models, and a small model TinyLlama-1.1B-Chat (Zhang et al., 2024a) as the expert model. Notice that our method only fine-tunes TinyLlama to possess the deep safety alignment property, following the method in Qi et al. (2025). Appendix C shows similar experimental results on more recent Llama3 models (Dubey et al., 2024).
@@
For utility, we consider two datasets: GSM8K (Cobbe et al., 2021) and Just-Eval (Lin et al., 2023). GSM8K is designed to test the model’s ability to solve complex mathematical problems. Notice that TinyLlama demonstrates very low accuracy on GSM8K and this task would challenging for the composite model. We use standard ROUGE-1 score for the accuracy on GSM8K, in line with existing evaluation practices. Similar to Xu et al. (2024), we use 1,000 diverse instructions from Just-Eval to evaluate the outputs in terms of helpfulness, clarity, factuality, depth, and engagement.
@@ 5.2 Experimental Results
**footnotetext: In Table 2, the Engaging score of Deep-Align on Vicuna is unusually high. We believe that this behavior is mainly due to the utility dataset used by Deep-Align fine-tuning, which influences Vicuna’s conversational style. See Appendix D for a detailed analysis.
SSD Avoids Over-refusal Responses on Borderline Queries.
As shown in Table˜3, SSD demonstrates a strong ability to handle borderline cases. Compared to the No Defense baseline, SSD introduces only a minimal increase in false refusals, indicating that its safety enhancement is well-targeted. In contrast, other defense methods exhibit higher FRRs, suggesting they are prone to over-refusal behavior. This result highlights that SSD effectively defends against genuine threats while maintaining helpfulness on nuanced, sensitive topics.
Method False Refusal Rate $\downarrow$
LLM as judge String matching
No Defense 20.00% 26.80%
Deep-Align 29.60% 49.20%
SafeDecoding 28.40% 65.60%
SSD (Ours) 20.80% 30.40%
@@
Table˜4 reports the Average Token Generation Time Ratio (ATGR), comparing the inference speed of our method, SSD, against other defense strategies like Deep-Align and SafeDecoding. The results validate that SSD can successfully accelerate decoding ($\text{ATGR}<1$), particularly for larger models. In contrast, SafeDecoding’s ATGR is greater than $1$ because its design of safety checks adds computational overhead and cannot utilize the auxiliary model’s outputs to accelerate inference. It is also worth noting that the ATGRs does not account for the cost of fine-tuning, which demands greater computational resources for large models.
@@ 7 Limitations
In this paper, the largest model on which we evaluate the proposed method is Llama2-13b. Due to limited GPU resources, we cannot conduct experiments to validate how our method applies to larger models like Llama2-70B. Another limitation of the current work is on the difference of conversational style between the original and the expert models, e.g., Vicuna and TinyLlama, which we do not investigate further. Future research can take into account this difference of conversational style to improve performance w.r.t. both model utility and safety for more practical scenarios.
@@ 8 Ethics Impact
This paper develops a lightweight and efficient method to strengthen existing LLMs with additional safety properties to defend against new jailbreak attacks. We empirically show that the developed method can effectively and efficiently equip the LLM with the deep safety alignment property and further improve the original model’s capacity against many types of jailbreaks. This research aims to enhance the safety of large models and to contribute positively to the broader field of AI research. We remark that the development of SSD only uses publicly available jailbreak prompts and do not create new ones. In the paper, only one jailbreak input query is exhibited in an abstracted way, for illustration purpose. We have also released our code to facilitate red-teaming efforts on LLMs.
Acknowledgment
This work is supported by the Innovation Project of Institute of Computing Technology, Chinese Academy of Sciences under Grant No. E561130 and the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDB0680101.
@@ B.4 Setup of Deep-Align
We train the Llama2-7b and Vicuna models with deep safety alignment properties using the same hyperparameters and datasets as the default settings in Qi et al. (2025). For TinyLlama used as the small expert model $m$, we adjust the learning rate of the hyperparameters to $2\times 10^{-4}$.
Appendix C Experimental Results on Llama3 Models
We conduct experiments on a more recent LLM Llama3.1-8b-Instruct (Dubey et al., 2024), to further validate the effectiveness of SSD. Here we use Llama3.2-1b-Instruct as the expert model because it shares the same vocabulary with Llama3.1-8b-Instruct. The results regarding safety and utility are presented in Table˜7 and Table˜8, respectively.
As shown in Table˜7, the original Llama3.1-8b model is highly vulnerable to prefilling attacks, with ASR of approximately 50%. Deep-Align provides a strong defense performance while the safety performance of SSD is close. This confirms that SSD effectively transfers the deep safety alignment property to this newer model. However, the utility evaluation in Table˜8 highlights the critical trade-offs between safety and utility. Our method shows much better utility performance compared with Deep-Align, particularly for the GSM8K score. Overall, SSD offers a comparable performance of the desired safety property while ensuring that the model’s utility remains largely intact.
Table 7: ASR (%) and harmful score for the Llama3.1-8b model under prefilling attack.
[…]
  10 tokens   20 tokens   40 tokens
Llama3.1-8b No Defense 46.73% (3.71) 49.84% (3.97) 50.16% (4.07)
Deep-Align 0.91% (1.08) 1.52% (1.14) 2.42% (1.16)
SSD 2.52% (1.59) 5.85% (1.69) 4.67% (1.71)
Table 8: Just-Eval and GSM8K scores of different defense methods on the Llama3.1-8b model.
Model Defense Just-Eval (1–5) $\uparrow$ GSM8K (%) $\uparrow$
Helpfulness Clarity Factuality Depth Engaging
Llama3.1-8b No Defense 4.23 4.89 4.80 3.97 4.35 81.8
Deep-Align 3.44 4.50 3.77 3.32 4.06 10.3
Ours 3.93 4.73 4.59 3.71 4.06 75.2
Appendix D Analysis of the Anomaly Engaging Score
In our main results in Table 2, we note that Deep-Align on Vicuna exhibits an anomalously high Engaging score of 4.36. We believe that the primary reason is that the anchor utility dataset used in Deep-Align promotes a more polite and conversational tone in Vicuna’s responses.
Table 9 provides a qualitative analysis comparing responses from both the original model (Baseline) and the Deep-Align model. Deep-Algin tends to introduce engaging openers like “Of course! I’m happy to help” or “Thank you for your question!”. To further validate this, we fine-tune Vicuna only on this utility dataset, which results in an even higher Engaging score of 4.46.
Another reason is the response length: the original model has an average response length of 179 words with Engaging score of 3.93, while the average length of Deep-Align is 227 words in Table 2. We then conduct a controlled-length experiment by constraining the output length via Prompting, and the results are reported in Table 10. We find that even with average length of 180 words, the Engaging score of Deep-Align is still much higher, suggesting that the response length is only a minor issue to the high Engaining score reported in Table 2.
Table 9: Comparison of response styles between the baseline Vicuna model and the Deep-Align model.
Query 1: What are some famous actors that started on Broadway?
Baseline There are many famous actors who got their start on Broadway. Here are a few examples…
Deep-Align Of course! I’m happy to help. Some famous actors who started their careers on Broadway include…
Query 2: How did US states get their names?
Baseline The United States is made up of 50 states…
Deep-Align Thank you for your question! The United States is a vast country…
Query 3: What is some cool music from the 1920s?
Baseline The 1920s was a decade of significant artistic and cultural change…
Deep-Align Of course! The 1920s were a time of great musical innovation…
Table 10: Engaging score of Deep-Align on Vicuna with different output length constraints. The undefended baseline model has an average length of 179 words and an Engaging score of 3.93.
Average Output Length 153 159 180 227
Engaging Score 4.24 4.29 4.31 4.42